OpenStudy (anonymous):

What is the solution of the system of equations 2x − 5y = 11 and −2x + 3y = −9? (1) (−3,−1) (3) (3,−1) (2) (−1,3) (4) (3,1)

OpenStudy (anonymous):

Show Work Please!!!

OpenStudy (anonymous):

were gonna do this together okay, real simple, first, we should add the first equation to the second equation okay: 2x − 5y = 11 −2x + 3y = −9 + --------------------- What do you get?

OpenStudy (anonymous):

-4x8y-2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

really? just look at this we have to add it like we would normally, it dosent matter if there are x and y's (2x)+(-2x)=0

OpenStudy (anonymous):

Fine I got the first step

OpenStudy (anonymous):

2x − 5y = 11 −2x + 3y = −9 + --------------------- -2y=2

OpenStudy (anonymous):

now can you find what y=?

OpenStudy (anonymous):

y = -1,3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what? how did you get y=-1,3, from -2y=2, you have to solve for y

OpenStudy (anonymous):

I don't know how to solve for Y

OpenStudy (anonymous):

if you have -2y=2, you dont know how to get y by ?itself

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

okay if i have -2y=2, then what should be done is divide both sidezs by -2, that gives us y=-1

OpenStudy (anonymous):

now, simply replace y=-1 into one of your orginal equation and get x

OpenStudy (anonymous):

What will be my final answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you have to solve this :2x − 5(-1) = 11

OpenStudy (anonymous):

Thanks for everything but I rather get the answer

OpenStudy (anonymous):

yeah, your gonna get far

OpenStudy (anonymous):

Thank You Sir.